### v1

# Coupled Flow-Thermal Analysis of a Rocket Nozzle with Charring Ablative Thermal Protection System

Abstract
This study investigates conjugate heat transfer analysis of a rocket nozzle featuring a charring ablative thermal protection system. The study is carried out by weakly coupling a commercial computational fluid dynamics (CFD) flow solver Fluent with an in-house material thermal response solver. The coupling of the two solvers is carried out by exchanging the boundary conditions at the fluid-solid interface using a non-iterative approach. Validation of the numerical framework is performed. A blowing correlation is used to model the pyrolysis gas behavior at the interface. Results highlight a significant surface heating and ablation particularly at the nozzle throat.
Nomenclature
$c_{p}$  = Specific heat capacity, J/kg K
$H_{\text{Pyro}}$  = Heat of pyrolysis, J/kg
$h$  = Convective heat transfer coefficient, W/m2 K
$\dot{m}_{c}^{\prime\prime}$  = Char ablation rate, kg/m2 s
$\dot{m}_{\text{Pyro}}^{\prime\prime}$  = Pyrolysis gas flow rate, kg/m2 s
$n$  = Normal coordinate
$P_{\text{eff}}$  = Effective thermophysical property of a material
$\dot{Q}_{\text{Pyro}}$  = Heat of pyrolysis per unit volume per unit time, W/m3
$\dot{Q}_{\text{Tr}}$  = Transpiration cooling rate per unit volume, W/m3
$Q$  = Heat of ablation (energy consumed per unit mass during ablation), J/kg
$\dot{q}^{\prime\prime}$  = Heat flux, W/m2
$S$  = Total recession thickness, mm
$\dot{s}$  = Surface recession rate, mm/s
$T$  = Temperature, K
$t$  = Time, s
$V$  = Volume, m3
Greek Symbols
$\epsilon$  = Surface emissivity
$\kappa$  = Thermal conductivity, W/m K
$\rho$  = Density, kg/m3
$\sigma$  = Stefan–Boltzmann constant, $5.67\times 10^{-8}$ W/m2 K4
Subscripts
blw  = Blowing corrected
c  = Char
conv  = Convective
g  = Gas
r  = Recovery
v  = Virgin
w  = Wall
$\infty$  = Freestream
1 Introduction
Rocket nozzle, a critical component of any propulsion system, must endure the extreme thermal loads during its operation. The high temperature and high speed gases exiting the combustion chamber interact with the nozzle wall and potentially degrade the nozzle material. In order to maintain and safeguard the structural integrity, a nozzle is incorporated with a thermal protection system (TPS). TPS materials, such as graphite, carbon-carbon, carbon-phenolic, silica-phenolic, and filled rubbers are particularly designed to sustain such extreme heat loads. These materials may undergo mass loss (ablation) when exposed to high-temperature environments in order to protect the underlying nozzle structure. The primary objective of a TPS is to shield the nozzle surface from the intense heat. Hence, the design of a TPS demands a comprehensive understanding of interactions between high temperature gas flow and TPS material. For the optimal performance of a rocket nozzle, it is essential to obtain an accurate thermal response of a TPS. While maximizing the TPS thickness may enhance the protection, but can put weight and space constraints. Conversely, minimizing the TPS thickness may compromise the protection, leading to system failure. A typical factor of safety used in the TPS design demands initial thickness to be at least twice the anticipated surface recession and 1.25 times the projected final char layer thickness [1, 2]. Therefore, designing a TPS that ensures the safety but maintaining minimal weight and space requirements is crucial.
Several studies to understand the behavior of gas flows through nozzles demonstrate that variations in nozzle geometry and operating conditions hugely influence its performance [3, 4, 5, 6, 7, 8, 9]. Research on nozzle surface ablation is broadly divided into two categories: experimental investigations [10, 11, 12] and computer simulations [13, 14, 15]. Experiments conducted by Sae-ngow et al.[11] analyzing ablation in static supersonic nozzles focused on silica phenolic composites. Their findings highlight that silica phenolic materials outperform rubber based thermal insulators under same operating conditions. Hui et al. [12] also performed experiments on a de Laval nozzle to investigate the relationship between nozzle geometry, gas flow behavior and ablation patterns. Their study revealed that the maximum ablation occurs at the throat of the nozzle. Despite experimental results providing valuable insights and being reliable, they are costly and time-consuming. These limitations have made researchers shift their focus on the computational studies.
Traditionally, thermal behavior and ablation analysis was decoupled, where both the flow solver and the material response solver were handled independently. Early methods like those proposed by Bartz [16] used simple correlations to calculate the convective heat flux. However, literature has shown that the material response of the nozzle is influenced by the surface temperature profile considered while calculating the wall heat flux [17]. As a result, such decoupled methods lacked fidelity to capture the intricate interactions between flow field and the solid domain.
A more accurate and robust approach to study the ablation of rocket nozzles is by employing conjugate heat transfer (CHT) analysis [18, 19, 20, 21, 22]. In such analysis, the flow solver is coupled with a material thermal response solver. The coupling of these solvers include sharing the boundary properties like heat flux and surface temperatures at the fluid-solid interface. Additionally, to account for the surface recession and the geometry changes due to ablation, the coupling between flow and thermal solver becomes essential [23].
The initial effort to perform a coupled analysis of flow field and ablation was done by NASA focusing on a pyrolyzing TPS material [24]. This study examined transient, one-dimensional ablation in the Stardust sample return capsule. However, the investigation did not account for surface recession due to ablation. Kuntz et al. at Sandia National Laboratories [23] performed a CHT analysis for the nose tip of IRV-2 vehicle, by integrating transient, two-dimensional ablation. However, this study had a limitation that they considered only non-charring TPS materials. Researchers at NASA researchers also analyzed the same test case using a similar approach [25], but their analysis was restricted to a one-dimensional ablation assumption.
Researchers at the University of Michigan [26, 27, 28] also conducted ablation analyses for IRV-2 vehicle and the Stardust reentry capsule using transient response from the material. However, their study was also limited to non-charring materials. Similarly, Thakre and Yang [29] investigated nozzle flows under the steady-state assumption focusing on non-charring materials with one-dimensional ablation. Cross and Boyd [18] performed ablation analysis for HIPPO nozzle, where they coupled the LeMANS solver and MOPAR-MD material solver. They used an algorithm exclusively developed and utilized by their team, which produced reliable ablation results. Since the methodology and code used were not made public, it is difficult for other researchers to advance and build upon their work.
The primary goal of the present study is to perform conjugate thermal analysis by loosely coupling a commercial CFD solver (ANSYS fluent) with an in-house material response code, charring ablator thermal solver (CATS). The surface temperature distribution obtained from the thermal response solver is transferred to the flow solver that gives surface heat flux to the thermal response solver, thereby enabling the interaction between the two solvers. A user defined function is utilized to apply the non-uniform temperature distribution on the nozzle wall in order to couple CFD solver with thermal solver. The study presents the temperature and heat flux distribution, and the resulting surface thermal ablation for a charring material.
The organization of the paper is as follows: the paper begins with a brief introduction of flow and material solvers, followed by a description of the flow-thermal coupling approach. A case from a technical NASA report [30] is used to validate the flow solver. After validation, the results for flow through a rocket nozzle are discussed with a primary focus on heat flux calculations and the resulting temperature rise within the TPS. Finally the paper concludes with a summary of key findings, highlighting the significance of the results.
2 Methodology
2.1 Governing Equations for Flow Solver [31]
For the nozzle flow problem, the governing equations consist of the continuity, momentum, and energy equations. Continuity equation ensures mass conservation through
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\vec{V})=0,$ (1)
where $\rho$ is density and $\vec{V}$ represents velocity.
For momentum conservation in each spatial direction $x,y,z$, the equation
$\frac{\partial(\rho\vec{V})}{\partial t}+\nabla\cdot(\rho\vec{V}\vec{V})=-\nabla p+\nabla\cdot\tau+\rho\vec{g}$ (2)
holds, incorporating pressure $p$, stress tensor $\tau$, and gravitational force $\vec{g}$.
Energy conservation is modeled through
$\frac{\partial(\rho E)}{\partial t}+\nabla\cdot\left(\vec{V}(\rho E+p)\right)=\nabla\cdot\left(k\nabla T\right)+\Phi,$ (3)
where $E$ denotes total energy, $T$ temperature, $k$ thermal conductivity, and $\Phi$ is the viscous dissipation term.
To accurately simulate turbulent flow in a nozzle, three turbulence models were utilized, viz., Spalart-Allmaras model, the Baseline $k$-$\omega$ model and the Shear Stress Transport (SST) $k$-$\omega$ model.
The Spalart-Allmaras (SA) model solves a single equation for turbulent viscosity:
$\frac{\partial(\rho\tilde{\nu})}{\partial t}+\nabla\cdot(\rho\tilde{\nu}\vec{V})=G_{\tilde{\nu}}-Y_{\tilde{\nu}}+S_{\tilde{\nu}}$, (4)
where $\tilde{\nu}$ represents modified turbulent viscosity.
The Baseline (BSL) $k$-$\omega$ model, combining the $k$-$\omega$ formulation for boundary layers with $k$-$\epsilon$ for the free stream, and and the Shear Stress Transport (SST) $k$-$\omega$ model, which blends boundary-layer sensitivity and free-stream behavior for better precision, solve transport equations for turbulent kinetic energy and specific dissipation:
$\frac{\partial(\rho k)}{\partial t}+\nabla\cdot(\rho k\vec{V})=P_{k}-\beta^{*}\rho k\omega+\nabla\cdot\left(\frac{\mu_{t}}{\sigma_{k}}\nabla k\right),$ (5)
$\frac{\partial(\rho\omega)}{\partial t}+\nabla\cdot(\rho\omega\vec{V})=\alpha\frac{\omega}{k}P_{k}-\beta\rho\omega^{2}+\nabla\cdot\left(\frac{\mu_{t}}{\sigma_{\omega}}\nabla\omega\right),$ (6)
Among these, the BSL $k$-$\omega$ model emerged as the most accurate for capturing boundary-layer separation and heat transfer in high-speed nozzle flows, a crucial aspect for accurate predictions.
2.2 Governing Equations for Material Response Solver
Charring ablative materials are widely used in TPS for high-speed re-entry vehicles, owing to their ability to dissipate heat by sacrificially losing mass through pyrolysis and ablation. The CATS [32] was utilized in this study to model this thermal response, accounting for heat conduction, pyrolysis, transpiration cooling, and surface recession.
The primary governing equation for heat conduction within charring ablative materials is expressed as
$\int_{V}\left(\rho c_{p}\frac{\partial T}{\partial t}\right)dV=\int_{A}\kappa(T)\frac{\partial T}{\partial n}dA-\int_{V}\left(\dot{Q}_{\text{pyro}}+\dot{Q}_{\text{trans}}\right)dV,$ (7)
where $T$ is the temperature, $\rho$ is the material density, $c_{p}$ is the specific heat capacity, $\kappa(T)$ is the temperature-dependent thermal conductivity, and $\dot{Q}_{\text{pyro}}$ and $\dot{Q}_{\text{trans}}$ represent the heat absorbed by pyrolysis and transpiration cooling, respectively. The thermal boundary conditions are defined by heat exchange at the exposed surface, where
$\kappa(T)\frac{\partial T}{\partial n}\bigg|_{\text{surface}}=\dot{q}_{\text{conv}}-\epsilon\sigma(T_{w}^{4}-T_{\infty}^{4})-\dot{m}_{c}Q^{*},$ (8)
with $\dot{q}_{\text{conv}}$ representing convective heat flux, $\epsilon$ emissivity, $\sigma$ the Stefan-Boltzmann constant, $T_{w}$ the wall temperature, $T_{\infty}$ the surrounding temperature, $\dot{m}_{c}$ the mass loss rate per unit area, and $Q^{*}$ the heat of ablation. The back surface is assumed insulated:
$\kappa(T)\frac{\partial T}{\partial n}\bigg|_{\text{back}}=0.$ (9)
The pyrolysis process, an endothermic reaction that decomposes material and releases pyrolysis gases, absorbs heat per unit volume as follows:
$\dot{Q}_{\text{pyro}}=\dot{\rho}H_{\text{pyro}},$ (10)
with $H_{\text{pyro}}$ denoting the heat of pyrolysis. Pyrolysis gases cool the char layer through transpiration cooling, modeled as
$\dot{Q}_{\text{trans}}=\dot{\rho}c_{p,g}(T-T_{g}),$ (11)
where $c_{p,g}$ is the specific heat of the gas and $T_{g}$ is the gas temperature. When the material surface reaches ablation temperature $T_{\text{abl}}$, surface recession begins, characterized by
$\dot{S}=\frac{\dot{m}_{c}}{\rho_{\text{char}}},$ (12)
with $\dot{S}$ as the recession rate, $\dot{m}_{c}$ the mass loss rate, and $\rho_{\text{char}}$ the char density.
Pyrolysis gas provides a major contribution in lowering the heat load on the surface of charring ablative TPS materials. This gas is a by-product of pyrolysis process and helps in heat removal from the material via transpiration cooling. It forms a thin protective layer between the incoming high temperature gases and the material. This study does not explicitly model the physical ejection of pyrolysis gases from the material surface. However, the impact of pyrolysis gas blowing on heat flux predictions is incorporated using a correlation function developed by Kays [33] and applied in recent research by Phadnis et al. [34]. The governing relationship is expressed as:
$\dot{q}^{\prime\prime}_{\text{blw}}=\Omega\dot{q}^{\prime\prime}_{\text{conv}},$ (13)
where the dimensionless blowing correction parameter $\Omega$ is calculated as:
$\Omega=\frac{\psi}{e^{\psi}-1},$ (14)
with $\psi$ being a non-dimensionalized function of the mass flux from the material into the boundary layer, given by:
$\psi=\frac{c_{p,g}\dot{m}^{\prime\prime}_{g}}{h},$ (15)
and the convective heat transfer coefficient $h$ is defined as:
$h=\frac{\dot{q}^{\prime\prime}_{\text{conv}}}{T_{r}-T_{w}}.$ (16)
Here, $\dot{q}^{\prime\prime}_{\text{conv}}$ is the convective heat flux, $T_{r}$ is the recovery boundary layer temperature, and $T_{w}$ is the wall temperature.
The blowing correction functions for heat flux, as expressed in Eqs. (13)–(16), are directly implemented in the thermal solver to account for the heat flux reduction due to pyrolysis gas blowing at the surface boundary. This approach ensures an accurate representation of the thermal response of the TPS material under extreme heat loads.
2.3 Flow-Thermal Coupling
In conjugate heat transfer of nozzle flows, the interaction between the hot gases and the material of the nozzle is referred to as flow-thermal coupling. In this work, it essentially involves the exchange of wall boundary conditions between the flow and solid domain of the nozzle. In order to achieve this, the commercial CFD flow solver, (ANSYS Fluent)[31], is coupled to an in-house material thermal response solver, Charring Ablator Thermal Solver (CATS). The coupling strategy enables the conjugate simulations of ablation of a charring and pyrolyzing material[35].
In literature, the flow-thermal solvers are coupled using different strategies. For instance [32] has coupled the flow and thermal solvers using two different ways: non-iterative and iterative coupled approaches. However this study was done for hypersonic reentry vehicles. The coupling strategy utilized in this study is based on non-iterative approach. This approach was seen to be computationally efficient as compared to the other method. In this method, the two solvers exchange the boundary conditions at the fluid-solid interface.
The interaction between flow and thermal response solver is outlined as follows:
  1. 1.
    A steady-state flow solution is obtained at the initial time point $t=t_{0}$, using ANSYS Fluent, with an isothermal boundary condition applied at the solid surface.
  2. 2.
    The solution of flow solver (convective heat flux distribution along the nozzle wall) is then passed to the in-house thermal response solver, CATS.
  3. 3.
    The thermal response code solves the transient heat conduction equation within the nozzle wall material, advancing in time from $t_{0}$ to $t_{0}+\Delta t$, with keeping the input heat flux constant. This captures the nozzle’s response over a small operational interval.
  4. 4.
    As the thermal response solver marches forward in time, at the new time point $t=t_{1}$, the updated wall temperature $T_{w}$ is transferred as a Dirichlet boundary condition to the flow solver.
  5. 5.
    Finally, using the updated surface temperature, the flow solver then computes a new steady-state convective heat flux distribution, thus providing input for the next thermal coupling step.
This entire process of coupling the properties at the fluid-solid interface relies on the coupling time-steps being very small. This ensures that the assumption of constant heat flux between two time steps is within tolerance limit.
3 Validation Case
In this study, ANSYS Fluent is utilized for the CFD simulations, while an in-house material thermal response solver is used for obtaining thermal response of a TPS material. The latter is already validated comprehensively in previous works [32, 35]. To validate the flow solver, ANSYS Fluent, a test case from the NASA Technical Report [30] is employed. The mesh of the flow domain is illustrated in Fig. 1. Details of the nozzle geometry are provided in Table 3. A structured mesh with 250 $\times$ 96 elements was used.
To accurately capture wall properties, the first layer of cells was positioned close to the wall to ensure adequate resolution of the boundary layer. Additionally, the mesh was refined near the nozzle wall to capture heat transfer phenomena effectively, maintaining the Y+ value below 1 throughout the nozzle wall. The test case parameters and boundary conditions are summarized in Table 3. An isothermal boundary condition was applied to the nozzle wall, with the wall temperature set to half of the nozzle inlet temperature.
Table 2: Nozzle Geometry Specifications
Parameter Value
Throat diameter 0.0458 m
Contraction area ratio (AR) 7.75:1
Expansion area ratio (AR) 2.68:1
Convergent half-angle 30°
Divergent half-angle 15°
Table 3: Operating Conditions
Condition Value
Nozzle inlet pressure $5.171\times 10^{5}$ N/m2
Nozzle inlet temperature 843.33 K
Nozzle inlet density 2.1306 kg/m3
Transport model Sutherland’s law
Turbulence model SA, BSL, SST
Exit Mach Number 2.5
Figure 1: Nozzle with cooling duct mesh
Figure 2: Heat transfer coefficient using different turbulence models
To validate the heat transfer coefficient variation along the wall, three different turbulence models are used: the Spalart-Allmaras (SA), Baseline (BSL) $k$-$\omega$, and Shear Stress Transport (SST) $k$-$\omega$ models. The results are provided in Fig. 2, wherein it can be noticed that the BSL and SA model show a good agreement with the experimental data.
The flow properties like mach number Fig. 3a, 3b, temperature Fig. 4a, 4b and pressure Fig. 5a, 5b with their respective axial variation can be visualized in the plots provided.
(a) Mach number contour
(b) Mach number variation along the axis
Figure 3: Mach Number Variation.
(a) Temperature contour
(b) Temperature variation along the axis
Figure 4: Temperature variation
(a) Pressure contour
(b) Pressure variation along the axis
Figure 5: Pressure variation.
4 Results And Discussion
After successful validation of flow solver, complete conjugate heat transfer analysis is performed. This time a thermal protection system (TPS) is introduced on the nozzle wall which is analyzed for the solid domain. TPS is made of AVCOAT 5026–39/HC-G, a glass-filled epoxy novolac ablative material. The thickness of 10 mm is provided for the solid domain simulations. AVCOAT material has an ablation temperature of 922 K and the relative thermo-physical properties of the material are taken from Ref. [36]. The mesh used for the flow simulations is same as used for validation case study, shown in Sec. 3. The mesh of the solid domain, as shown in Fig. 6, is having 250 cells in transverse direction and 50 cells in radial direction. The simulation results are provided in different time intervals in order to give a clear picture of results.
Figure 6: Mesh for the thermal protection system
4.1 Comparison of Heat Flux and Surface Temperature
Figure 7a shows the convective heat flux variation along the nozzle during the first second of operation. The same heat flux is applied to the TPS material to evaluate the thermal response of the AVCOAT material. It can be observed that the maximum heat flux occurs in the throat region of the nozzle. This corresponds to the maximum surface temperature in this region, as can be seen from Fig. 7b. Additionally, it can be observed that the heat flux shows a decreasing trend with increasing time in this time interval.
(a) Heat flux variation
(b) Surface temperature variation
Figure 7: Comparison of surface heat flux and surface temperature; Time interval: 0.2-1 s.
Figure 8a depicts heat flux variation during the time interval of one to five seconds. It can be observed that the heat flux generally shows a decreasing trend over time. However, during the 5th second, a reversal in this trend is noted. As seen in Fig. 8b, it can be noticed that the surface temperature of the TPS material near throat region reached the ablation temperature during the 5th second. The rise of surface temperature to this limit shows the material in the throat region goes under ablation earlier as compared to other regions.
(a) Heat flux variation
(b) Surface temperature variation
Figure 8: Comparison of surface heat flux and surface temperature; Time interval: 1-5 s.
During the time interval of 5 to 10 s, the heat flux showed a complete reversal in heat flux trend, as depicted in Fig. 9a. A gradual and consistent increment in the heat flux is evident over time. From the temperature profile, Fig. 9b, it can be seen that a large portion of the nozzle wall has reached the ablation threshold of the material. Furthermore, the results illustrate that the ablation is spreading to the adjacent regions of the throat area as specified by the expansion of both heat flux and temperature profiles.
(a) Heat flux variation
(b) Surface temperature variation
Figure 9: Comparison of heat flux and surface temperature; Time interval: 5-10 s.
Between 10 and 15 s, the heat flux stabilizes across the nozzle wall. It becomes apparent that the increment in the heat flux is gradual, indicating it is approaching a steady state, as shown in Fig. 10a. From the temperature profile (Fig. 10b), it is clear that the most part of nozzle wall has undergone ablation during this time interval.
(a) Heat flux variation
(b) Surface temperature variation
Figure 10: Comparison of heat flux and surface temperature; Time interval: 10-15 s.
During the 15-20 s interval, the progression in the heat flux is minimal, as illustrated in Fig. 11a. The surface temperature of the nozzle wall, with the exception of entry section reached the ablation threshold of 922 K, except the entry part of the nozzle, as highlighted in Fig. 11b
(a) Heat flux variation
(b) Surface temperature variation
Figure 11: Comparison of heat flux and surface temperature; Time interval: 15-20 s.
As presented in Fig. 12a, the heat flux in the interval of 20 to 25 s reaches a stable state. There is no further increase in the heat flux over time. From the Fig. 12a, it is clear that the heat flux at 22 s and 25 s is identical. The surface temperature of the nozzle wall reaches an isothermal temperature 922 K in this time, as observed in Fig. 12b.
Simulation results were extended further in time, up to 120 s. Since the nozzle wall temperature remained at 922 K, no further increment in the heat flux was evident. The heat flux at 120 s matched completely with that at 22 s. At this point, the entire nozzle wall has undergone ablation. The temperature contour of the TPS material is presented in Fig. 4a.
(a) Heat flux variation
(b) Surface temperature variation
Figure 12: Comparison of heat flux and surface temperature; Time interval: 20-25 s
4.2 Throat Region Heat Transfer Analysis
The throat of a rocket nozzle is a critical region as it separates subsonic flow region from the supersonic flow, hence, a detailed analysis is needed. This region of the nozzle experiences maximum heat flux, as previously discussed. The results of heat flux variation with time for the throat region are highlighted in Fig. 13a. The results illustrate that, initially the heat flux is at its highest value, followed with a sudden decrease in the heat flux as time progresses. As demonstrated by the graph, it is clear that the ablation process begins at around 5 s. Once ablation starts, heat flux increases again for a few seconds. Further in time, one can see that heat flux attains a stable value before stabilizing at a constant value.
(a) Nozzle throat heat flux variation
(b) Nozzle throat temperature variation
Figure 13: Comparison of heat flux and temperature at nozzle throat
Figure 14: Comparison of heat flux at three different X locations
In comparison to the heat flux, the surface temperature of the throat region continues to increase over time, until it reaches the ablation temperature of 922 K, as shown in Fig. 13b. The temperature remains constant over time, once the surface of the throat reaches the ablation temperature. Simulation results are obtained up to 120 s, but results are shown only up to 60 s, since no further changes in the heat flux are observed beyond this time.
4.3 Comparative Heat Flux Analysis Across Nozzle Regions
To further understand the progression of heat transfer within the TPS, three locations are investigated, viz., the throat (X = 0.091 m), the converging section (X = 0.075 m), and the diverging section (X = 0.105 m) of the nozzle.The heat flux results are presented in Fig. 14. All three locations follow a similar trend, as discussed for the throat region. It can be noted that the maximum heat flux is noticed in the throat region, followed by the diverging section, with the converging section experiencing the least heat flux.
The temperature variation through the thickness of the TPS material at these locations is also analyzed, as depicted respectively in Figs. 15b, 15a and 16b. The surface near the high temperature gas flow reaches a notably higher temperature, while the back-end of the material remains at the adiabatic wall temperature. In addition to that, a gradual decrease in temperature through the thickness of the TPS material is observed.
(a) Temperature through the material at X= 0.091 m (throat)
(b) Temperature through the material at X= 0.075 m
Figure 15: Temperature variation at different locations.
(a) Temperature through the material at X= 0.105 m
(b) Temperature contour of TPS
Figure 16: Temperature through the material at X= 0.105 m and temperature contour of TPS.
4.4 Surface Recession
Once the surface temperature of the nozzle TPS reaches the ablation temperature, the surface of the nozzle wall begins to erode due to surface recession. The surface recession over time is calculated by the Eq. (12). The study of surface recession is critical as the dynamics of nozzle flow can notably get affected due to the surface ablation.
The initial and the final nozzle wall geometries are shown in Fig. 17a. The receded geometry is plotted at 120 s of nozzle operation. It is evident that the maximum surface recession occurs at the throat of the nozzle compared to the other regions of the nozzle wall. The maximum recession at the nozzle throat is calculated to be approximately 2.5 mm, as can be visualized in Fig. 17b. The surface recession of TPS material aligns closely with the heat flux and surface temperature variations.
(a) Initial vs receded nozzle wall
(b) Total surface recession
Figure 17: Surface recession.
5 Conclusions
A conjugate heat transfer analysis of rocket nozzle equipped with charring ablative thermal protection system was evaluated under the extreme heat loads. The effect of pyrolysis gas at the fluid and solid interface was modeled using blowing correction correlation. The coupled study of flow (steady state analysis) and transient thermal solver CATS enabled a detailed analysis of TPS of a rocket nozzle. The validation of the flow solver with the established experimental data demonstrated the reliability of our flow simulations. The thermal response modeled by in-house solver helped to understand the mechanisms, such as, ablation, charring and pyrolysis within the TPS material. Surface recession was calculated for 120 s, illustrating how the nozzle surface erodes away under extreme gas temperatures.
The maximum surface recession occurred in the throat region with a magnitude of approximately 2.5 mm at 120 s. These findings of ablation and surface recession underscore the need for enhanced heat insulation in such extreme heat load regions. The used approach presents a framework for predicting the thermal behavior of AVCOAT TPS material under extreme conditions. Nevertheless, there are areas for improvement. Future studies will be performed to understand the impact of coupling time-step better. Furthermore, this study did not incorporate mesh updates for the flow domain; incorporating mesh adaptation in future studies could capture the dynamics of flow better. This would enable even more accurate predictions of TPS performance and durability.
Acknowledgments
The authors acknowledge the National Supercomputing Mission (NSM) for providing computing resources of “PARAM Sanganak” at IIT Kanpur, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and the Department of Science and Technology (DST), Government of India.
Author Declarations
The authors have no conflict of interest to disclose.
Data Availability
The data that support the findings of this study can be made available by the corresponding author upon a reasonable request.